y'' − y' + 1/4 y = 8 + ex/2

1
Expert's answer
2021-05-11T06:33:08-0400

y'' − y' + 1/4 y = 8 + ex/2


"= y''(x)-y'(x)+\frac{1}{4}y(x)=8+e^{x\over2}"


the general solution will be sum of the complementary solution and particular solution.


we find the complementary solution by solving "\frac{d^2y(x)}{dx^2}-\frac{dy(x)}{dx}+\frac{y(x)}{4}=0"


comes out to be

"y(x)=y_1(x)+y_2(x)"


"y(x)=c_1e^{x\over 2}+c_2e^{x\over 2}" .........(1)


now find particular solution

"\frac{d^2y(x)}{dx^2}-\frac{dy(x)}{dx}+\frac{y(x)}{4}=e^{x\over 2}"


comes out to be

"y_p(x)=\frac{1}{2}e^{x\over 2}x^2+32" ............(2)


combining both (1) and (2) we get the answer


"\boxed{y(x)=c_2e^{x\over 2}x+c_1e^{x\over 2}+\frac{1}{2}e^{x\over 2}x^2+32}answer"